OpenStudy (anonymous):

Could I have some assistance please? - if you're driving 24 km in a car and you're driving 60 km/hr how fast will you get there?

OpenStudy (anonymous):

Just divide it

OpenStudy (anonymous):

what do youmean @iambatman like 60/24?

OpenStudy (anonymous):

You're looking for how much time it would take right to travel 24 km when going 60 km/h?

OpenStudy (anonymous):

just imagine that 1 hr you drive 60km but you ony have 24km to drive . it will be 1min 1km so 24km in 24min

OpenStudy (anonymous):

Divide 24/60 = 0.4 * 60 min = 24 minutes.

OpenStudy (anonymous):

you divided and multiplied by 60 minutes (hour) and that was the result?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

is that the correct format , so i can use on other simular problems?

OpenStudy (anonymous):

and 24 is my answer o.o? @iambatman wow im dumb.

OpenStudy (anonymous):

Pay attention to the units and it's easy. :)

OpenStudy (anonymous):

Yes your answer is 24 minutes.

OpenStudy (anonymous):

i feel like a sped wow, i need to work on my paying attention lol.

OpenStudy (kc_kennylau):

Speed is defined to be distance over time: s=d/t And we can derive to more formulas from that: d=st, t=d/s

OpenStudy (kc_kennylau):

Memorize these 3 formulas by heart and you'll be able to do them quickly in no time

OpenStudy (anonymous):

Good ol physics! Don't know why I didn't mention that.

OpenStudy (kc_kennylau):

s=d/t, d=st, t=d/s

OpenStudy (anonymous):

You guys are so helpful :')

OpenStudy (anonymous):

4 am..I should've told you that haha, thank you @kc_kennylau

OpenStudy (anonymous):

its 6am here :D we dont sleep

OpenStudy (kc_kennylau):

6pm here :DDDDDD

OpenStudy (fibonaccichick666):

6:17... just had to chime in. fun read guys

OpenStudy (fibonaccichick666):

am

OpenStudy (anonymous):

d=v/t is actually a very powerful formula. You can figure out many things from it, such as how much time does it take light to reach us from the sun, given speed of light and distance from the sun. Distance of the sun from earth: 1.496*10^11 m Speed of light ~ 3.0 *10^8 m/s v=d/t => t =d/v t = (1.496*10^11 m)/ (3*10^8 m/s) = 498.6s = 8.31 minutes. Sorry for the randomness, just wanted to show something related to physics. Haha.